# LinkPrompt: Natural and Universal Adversarial Attacks on Prompt-based Language Models

### v1

@@ C.2 Prompt used for fine-tuning Llama2
Predict the "[mask]" with "politics", "sports", "business" or "technology" to make the whole sentence semantically natural: Najaf battle a crucial test for Allawi Clashes between US troops and Sadr militiamen escalated Thursday, as the US surrounded Najaf for possible siege. [mask] news => politics; Galaxy, Crew Play to 0-0 Tie (AP) AP - Kevin Hartman made seven saves for Los Angeles, and Jon Busch had two saves for Columbus as the Galaxy and Crew played to a 0-0 tie Saturday night. [mask] news => sports; Wall St. Bears Claw Back Into the Black (Reuters) Reuters - Short-sellers, Wall Street’s dwindling band of ultra-cynics, are seeing green again. [mask] news => business; Oracle expands midmarket ambitions Company looks to juice its application server business with a version tuned for smaller organizations. [mask] news => technology
SST2
@@
Predict the "[mask]" with "bad" or "good" to make the whole sentence semantically natural: a stirring , funny and finally transporting re-imagining of beauty and the beast and 1930s horror films It was [mask]. => good; apparently reassembled from the cutting-room floor of any given daytime soap .It was [mask]. => bad
IMDB
@@
Predict the "[mask]" with "bad" or "good" to make the whole sentence semantically natural: Not only is it a disgustingly made low-budget bad-acted movie, but the plot itself is just STUPID!!!<br /><br />A mystic man that eats women? (And by the looks, not virgin ones)<br /><br />Ridiculous!!! If you´ve got nothing better to do (like sleeping) you should watch this. Yeah right. It was [mask]. => bad; Went to see this as Me and my Lady had little else to do on a sunday afternoon I like films that deal with sleazy,loser characters and this is full of em. After a slow start we get some good turns from the cast but it is the actual ’Bellini’ that both makes and lets the film down. The ’Bellini’ is one of the funniest scenes I have seen in a film for a long while but is too short and could have made this a masterpiece overall 71/2 out of 10 It was [mask]. => good
HATE
@@
Predict the "[mask]" with "harmless" or "hate" to make the whole sentence semantically natural: Happy birthday to my brother @DavonteJones10 hope you have a good day fam love you &#128170;&#127998; see you later bro [mask] speech => harmless; RT @bateson87: Send Barkley off. He’s a dirty bastard [mask] speech => hate
FN
@@
Predict the "[mask]" with "real" or "fake" to make the whole sentence semantically natural: There was better news today for the ex-Toon hero Kevin Keegan when, after having resigned from the Newcastle United manager’s job last week, he was offered a new job: Holding Joey Barton’s Coat. Barton, still a player at United, albeit with a six-match ban, is a rabble-rouser, a trouble causer, a bit ’handy’, pushy, a ’lad’, good with his fists… temperamental, know what I mean? He regularly gets into fights, and is always in need of someone to ’ hold his coat’. The last time Barton got into a ’scuffle’, it ended in a jail sentence, and prior to that, he left Man City teammate Ousmane Dabo with his face ’caved in’. On both occasions, he was wearing a jacket, and believes he could have done so much more damage had he had a ’second’ to hold his apparel for him. Ex-boss Keegan regarded himself as something of a father figure to Barton at Newcastle, defended him in front of the Newcastle board, and stood by him when he emerged from the nick recently. Now King Kev is to support the lad permanently as he follows him around, waiting for him to explode. The job is Full Time, 24 hours a day, 7 days a week, 52 weeks a year. You get the idea. Said Keegan:" I’m lookin forward to it. He’s a nice lad, wears his heart on his sleeve, but there’s nuthin wrong with that. He’s got a fiery temperament, but he gives it his all, and you can’t ask for anythin more than that." It was [mask]. => real
FR
@@
Predict the "[mask]" with "real" or "fake" to make the whole sentence semantically natural: This is a great product. No more fears of loosing food to bears. No assaults yet but expect it to hold up nicely. [mask] review => real; Best FPV training transition to FPV and the FPV class is a lot of fun! The other two FPV classes are a bit more complex [mask] review => fake

### v2

@@ C.1 Prompt used for naturalness measurement via ChatGPT
You are responsible for deciding which expression is more natural. Just respond with a number. Sentence1: {sen1}; Sentence2: {sen2}.
@@ C.2 Prompt used for fine-tuning Llama2
AG Predict the "[mask]" with "politics", "sports", "business" or "technology" to make the whole sentence semantically natural: Najaf battle a crucial test for Allawi Clashes between US troops and Sadr militiamen escalated Thursday, as the US surrounded Najaf for possible siege. [mask] news => politics; Galaxy, Crew Play to 0-0 Tie (AP) AP - Kevin Hartman made seven saves for Los Angeles, and Jon Busch had two saves for Columbus as the Galaxy and Crew played to a 0-0 tie Saturday night. [mask] news => sports; Wall St. Bears Claw Back Into the Black (Reuters) Reuters - Short-sellers, Wall Street’s dwindling band of ultra-cynics, are seeing green again. [mask] news => business; Oracle expands midmarket ambitions Company looks to juice its application server business with a version tuned for smaller organizations. [mask] news => technology
@@
SST2 Predict the "[mask]" with "bad" or "good" to make the whole sentence semantically natural: a stirring , funny and finally transporting re-imagining of beauty and the beast and 1930s horror films It was [mask]. => good; apparently reassembled from the cutting-room floor of any given daytime soap .It was [mask]. => bad
@@
IMDB Predict the "[mask]" with "bad" or "good" to make the whole sentence semantically natural: Not only is it a disgustingly made low-budget bad-acted movie, but the plot itself is just STUPID!!!<br /><br />A mystic man that eats women? (And by the looks, not virgin ones)<br /><br />Ridiculous!!! If you´ve got nothing better to do (like sleeping) you should watch this. Yeah right. It was [mask]. => bad; Went to see this as Me and my Lady had little else to do on a sunday afternoon I like films that deal with sleazy,loser characters and this is full of em. After a slow start we get some good turns from the cast but it is the actual ’Bellini’ that both makes and lets the film down. The ’Bellini’ is one of the funniest scenes I have seen in a film for a long while but is too short and could have made this a masterpiece overall 71/2 out of 10 It was [mask]. => good
@@
HATE Predict the "[mask]" with "harmless" or "hate" to make the whole sentence semantically natural: Happy birthday to my brother @DavonteJones10 hope you have a good day fam love you &#128170;&#127998; see you later bro [mask] speech => harmless; RT @bateson87: Send Barkley off. He’s a dirty bastard [mask] speech => hate
@@
FN Predict the "[mask]" with "real" or "fake" to make the whole sentence semantically natural: There was better news today for the ex-Toon hero Kevin Keegan when, after having resigned from the Newcastle United manager’s job last week, he was offered a new job: Holding Joey Barton’s Coat. Barton, still a player at United, albeit with a six-match ban, is a rabble-rouser, a trouble causer, a bit ’handy’, pushy, a ’lad’, good with his fists… temperamental, know what I mean? He regularly gets into fights, and is always in need of someone to ’ hold his coat’. The last time Barton got into a ’scuffle’, it ended in a jail sentence, and prior to that, he left Man City teammate Ousmane Dabo with his face ’caved in’. On both occasions, he was wearing a jacket, and believes he could have done so much more damage had he had a ’second’ to hold his apparel for him. Ex-boss Keegan regarded himself as something of a father figure to Barton at Newcastle, defended him in front of the Newcastle board, and stood by him when he emerged from the nick recently. Now King Kev is to support the lad permanently as he follows him around, waiting for him to explode. The job is Full Time, 24 hours a day, 7 days a week, 52 weeks a year. You get the idea. Said Keegan:" I’m lookin forward to it. He’s a nice lad, wears his heart on his sleeve, but there’s nuthin wrong with that. He’s got a fiery temperament, but he gives it his all, and you can’t ask for anythin more than that." It was [mask]. => real
@@
FR Predict the "[mask]" with "real" or "fake" to make the whole sentence semantically natural: This is a great product. No more fears of loosing food to bears. No assaults yet but expect it to hold up nicely. [mask] review => real; Best FPV training transition to FPV and the FPV class is a lot of fun! The other two FPV classes are a bit more complex [mask] review => fake